y=x-4 (I know the answer would be 4,-4, but I'm not sure why that would be correct. Please help
Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

y=x-4 (I know the answer would be 4,-4,
but I'm not sure why that would be correct.
----
(4,-4) is NOT correct.
----
If x = 4, y = 4-4 = 0
(4,0) is correct.
======================
Cheers,
Stan H.

Answer by Alan3354(69443)

y=x-4 (I know the answer would be 4,-4, but I'm not sure why that would be correct.
-------------
y = x - 4 has an infinite # of "answers"
Pick a number for x and find y.
-----------
As for (4,-4), that's not one of them
If x = 4, then y = 4-4 = 0 --> (4,0)
--------------
If you graph it, you get a straight line, and every point on it is a solution, or answer.